# Roundtable Policy: Improving Scientific Reasoning and Narratives through Confidence-Weighted Consensus of LLMs

### v1

@@ Abstract
Large language models (LLMs) have demonstrated remarkable capabilities not only in language generation but also in advancing scientific discovery. A growing body of work has explored ways to improve their reasoning, from self-consistency and chain-of-thought to multi-agent debate. Inspired by the dynamics of scientific committees and The Society of Mind, we introduce Roundtable Policy, a complementary inference-time reasoning framework that performs inference through the weighted consensus of multiple LLMs. Our findings indicate that this approach significantly enhances reasoning in complex heterogeneous scientific tasks and improves scientific narratives in terms of creativity, rigor, and logical coherence, while reducing hallucinations that single models are prone to. Our approach emphasizes structured and interpretable consensus rather than opaque convergence, while requires only black-box access and uniform procedures, making it broadly applicable to multi-LLM reasoning. Project website at https://ai4scidisc.github.io/roundtable/.
@@ 1 Introduction
Large language models (LLMs) have demonstrated impressive capabilities in text generation, reasoning, and few-shot adaptation. These advances have opened exciting opportunities for scientific discovery, where complex, interdisciplinary problems demand not only accurate reasoning but also coherent narrative construction. However, current LLMs may hallucinate facts, introduce implausible logical jumps, or produce fragmented explanations. Such limitations are particularly problematic in scientific contexts, where reliability and clarity are equally critical. For instance, a scientific proposal must not only provide factually correct content, but also organize it into a structured and persuasive narrative. Addressing these dual requirements calls for mechanisms that go beyond the capacity of a single LLM.
@@
A natural motivation is therefore to leverage multiple LLMs, aggregating their complementary strengths. Prior work has explored self-consistency with multiple sampled reasoning paths (Wang et al., 2022), multi-agent debate frameworks (Du & et al., 2023; Madaan & et al., 2023), and tool-augmented multi-agent systems (Schick et al., 2023; Ghafarollahi & Buehler, 2024). However, these approaches are often limited to single-task reasoning, rely on emergent dynamics that are difficult to interpret, or lack systematic mechanisms for weighting contributions across heterogeneous domains. Ensemble-style methods such as majority voting treat all models equally, overlooking that different models excel in different domains or evaluation dimensions. Debate-style methods, while sometimes converging to better outputs, remain inherently opaque: it is often unclear why a particular answer was selected, or hard to quantify which models were more trustworthy in the process. In scientific applications, this lack of transparency poses a major barrier, since researchers require not only correct outputs but also interpretable evidence supporting the decision.
At a broader level, scientific problem solving is inherently multi-faceted, requiring both precise reasoning and coherent long-form narratives. While individual LLMs have made impressive progress, they still face challenges in consistently meeting these diverse demands across domains and tasks. This motivates the need for a structured approach to harness collective intelligence—one that integrates complementary expertise, surfaces and resolves disagreements transparently, and preserves accountability in the decision process.
Building on these considerations, we introduce Roundtable Policy, an inference-time framework that formalizes deliberation among multiple LLMs into a principled process for trustworthy scientific reasoning and narrative generation (Figure 1). In our approach, diverse models first propose candidate responses, which are then jointly evaluated to record their performances. These evaluations are distilled into a confidence-weight table that encodes both historical reliability and uncertainty for each candidate across tasks and dimensions. At inference, this confidence-weight table guides a fusion process that integrates the diverse responses into a single consensus answer—balancing factual accuracy with narrative coherence while preserving transparency in how the decision is reached. As illustrated in Figure 1, even on a scientific question such as explaining why nucleation accelerates at greater undercooling while growth dominates at smaller undercooling, individual models provide partial or inconsistent accounts. The roundtable process fuses these perspectives into a coherent explanation, showing how consensus-driven reasoning can elevate fragmented insights into a more complete scientific narrative.
We evaluate Roundtable Policy in two complementary scientific regimes. In the MultiTask setting, heterogeneous problems from geoscience, biology, mathematics, and materials science are solved simultaneously, testing cross-domain reasoning. In the SingleTask setting, models generate structured scientific proposals, testing long-form narrative construction. Across both, our findings show that Roundtable Policy consistently improves factual reliability, reasoning quality, and narrative coherence, yielding an average 13.01% improvement in MultiTask and 11.04% improvement in SingleTask compared to individual models.
Unlike ensemble-style methods such as majority voting, which treat all models equally, our approach adaptively weights models based on demonstrated reliability. And in contrast to debate-style frameworks, whose convergence dynamics are opaque, Roundtable Policy produces explicit, auditable signals that make the aggregation process transparent. This combination of principled weighting and interpretability distinguishes our method as a scalable and accountable alternative for multi-model scientific reasoning.
In summary, our contributions are threefold: (i) we propose Roundtable Policy, an inference-time framework for performance-weighted consensus aggregation among multiple LLMs; (ii) we demonstrate that this approach enhances both factual reasoning and scientific narrative across heterogeneous tasks; and (iii) we release a graduate-level materials science benchmark derived from UCLA PhD qualifying exams, which serves as a rigorous, domain-authentic testbed and complements existing evaluations of scientific LLMs. More broadly, the principles underlying Roundtable Policy apply across multi-model scientific workflows, offering an interpretable and generalizable path toward AI-assisted reasoning and communication in science.
@@ 2 Methodology
This section formalizes the Roundtable Policy framework. We begin with the intuition behind the Roundtable Policy and then present the task settings and notation (a complete list of symbols is provided in Appendix §A). We next describe how the confidence-weight table is learned from repeated interactions between players and graders, and how it subsequently guides inference through consensus-weighted fusion.
@@ 2.1 Design Intuitions behind Roundtable Policy
Scientific decision-making rarely relies on a single voice. In research committees, for example, each member independently evaluates a proposal, but the final judgment is not determined by a simple vote. Instead, reviewers’ assessments are weighed by domain expertise, reliability, and the clarity of justification before being synthesized into a collective decision. This process balances agreement with accountability, ensuring that the outcome is both credible and transparent.
Roundtable Policy draws on the same principle. Rather than treating LLMs as isolated predictors, we view them as members of a committee. A set of players propose candidate responses informed by diverse inductive biases. These are evaluated by an ensemble of graders, whose rubric-based assessments provide explicit signals of quality and uncertainty. Finally, a structured fusion step aggregates these evaluations into a single consensus output—analogous to how a committee chair synthesizes diverse reviews into a reasoned decision.
This design is further motivated by the perspective of The Society of Mind (Minsky, 1988), which frames intelligence as the coordination of many specialized, relatively simple agents. By embracing this philosophy, Roundtable Policy transforms the inherent diversity of LLMs into a strength, enabling collective reasoning that is both interpretable and robust.
2.2 Task Settings
We evaluate Roundtable Policy in two complementary scientific regimes that mirror the dual demands of scientific reasoning: solving pointwise problems and constructing coherent narratives.
MultiTask. Models are simultaneously queried with nine heterogeneous scientific questions spanning geoscience, biology, mathematics, and materials science (with six distinct sub-categories). Task formats include multiple-choice, fill-in-the-blank, and short-answer questions. Intuitively, this setting resembles a written exam with diverse problem types—success requires breadth of expertise rather than mastery of a single domain.
SingleTask. Models are prompted to generate structured scientific proposals following a standardized three-paragraph template with sections for Background, Methodology, and Impact. This setting captures the narrative aspect of scientific work: beyond producing isolated facts, one must weave them into a logically consistent and persuasive storyline.
Across both regimes, we deploy pretrained LLMs (players) to generate candidate responses and LLMs (graders) to form the AI Committee. Just as in a real committee, players provide diverse answers while graders act as evaluators who judge each contribution against shared standards. In MultiTask, graders jointly evaluate the responses per round, analogous to examiners scoring a large pool of student answers. In SingleTask, graders assess proposals paragraph-by-paragraph along three rubric dimensions—creativity, scientific rigor, and logical coherence—reflecting the multifaceted criteria by which scientific proposals are judged.
@@ 2.2 Task Settings
Each evaluation yields a trust weight with two components: (i) a quality score in , where negative values indicate incorrect or misleading responses, and (ii) a calibrated 95% confidence interval representing grader certainty. Together, these form the equivalent of both a “grade” and a measure of “grading confidence,” ensuring that downstream decisions reflect not only evaluation outcomes but also their reliability. Notation follows Appendix §A denotes the -th player agent, the -th grader, and the fusion agent. Player responses and grader scores are indexed by round .
2.3 Learning the Confidence-Weight Table
The confidence-weight table serves as a structured memory of each player’s historical reliability across subtasks (MultiTask) or rubric dimensions (SingleTask). Unlike fine-tuning approaches, Roundtable Policy does not alter the parameters of the underlying LLMs. Instead, is iteratively updated from repeated rounds of evaluation, much like how a committee gradually learns which members are most reliable in which topics.
@@ 2.3 Learning the Confidence-Weight Table
At round , a query is sampled from the task dataset. Players generate responses , which are pooled into . Graders then evaluate these responses, producing , which are averaged across graders to yield . After rounds, the cumulative evaluations are distilled into:
@@ 2.4 Inference with the Confidence-Weight Table
Once has been learned, it guides consensus fusion at inference. Given a new query , each player generates a response . The collection is passed to the fusion process , which aggregates them using the pretrained :
Here, ensures that expertise is applied where it matters most. In MultiTask, it emphasizes domain-specific reliability (e.g., assigning higher weights to players consistently strong in mathematics). In SingleTask, it captures stylistic and analytical strengths across rubric dimensions, reflecting how different voices contribute to different aspects of a proposal.
The fusion process itself is not fine-tuned; it acts as an orchestrator that conditions on to produce a consensus answer. This mimics how a committee chair integrates both the content and credibility of reviewers before issuing a final recommendation. Optionally, the AI Committee can re-evaluate to assign a quality score and confidence interval, preserving interpretability and accountability throughout the process.
@@ 3 Experiments
We design experiments to systematically evaluate the effectiveness and robustness of the proposed Roundtable Policy. Our analysis is guided by the following research questions:
RQ1: To what extent does Roundtable Policy enhance scientific reasoning and writing performance relative to individual LLMs? (see §3.1)
RQ2: How stable and robust is Roundtable Policy over long-run evaluation regimes, and does it consistently provide more reliable performance? (see §3.2)
RQ3: To what degree do AI graders introduce bias, and can the AI Committee mitigate these biases to achieve greater inter-grader consistency? (see §3.3)
3.1 Performance Improvement (RQ1)
| Setting | Model | Material Science Dataset | Others | |||||||
@@ 3.1 Performance Improvement (RQ1)
MultiTask |
GPT-4o | 85.894.27 | 78.265.37 | 87.714.11 | 74.595.49 | 85.144.40 | 85.584.60 | 87.812.13 | 83.552.20 | 51.536.64 |
| Grok-2 | 83.924.07 | 76.244.96 | 85.613.92 | 76.184.99 | 82.024.10 | 82.634.39 | 88.841.70 | 77.392.28 | 51.316.25 | |
| Claude-3.5H | 82.074.65 | 71.206.04 | 84.724.37 | 71.495.83 | 77.745.03 | 77.955.22 | 81.192.58 | 73.682.58 | 48.526.75 | |
| GPT-4.5p | 72.866.77 | 71.466.96 | 76.676.12 | 72.786.36 | 77.185.71 | 77.686.17 | 82.502.68 | 82.592.08 | 50.636.89 | |
| DeepSeek-chat | 77.635.35 | 72.026.14 | 81.534.97 | 72.965.79 | 79.644.87 | 79.815.23 | 72.233.16 | 78.052.41 | 47.836.57 | |
| Claude-3.7S | 77.433.46 | 70.514.05 | 79.243.13 | 72.393.70 | 77.413.49 | 79.493.12 | 62.342.93 | 72.561.97 | 47.765.36 | |
| Llama-3.3 | 72.145.96 | 63.866.95 | 75.335.55 | 67.136.30 | 71.235.71 | 73.875.72 | 87.431.93 | 72.772.49 | 48.646.54 | |
| Qwen-Plus | 80.325.01 | 72.366.03 | 82.974.71 | 73.425.66 | 81.494.70 | 81.184.93 | 86.992.08 | 76.272.48 | 51.646.57 | |
| Gemini-2.0 | 67.267.61 | 61.618.09 | 74.966.58 | 65.737.31 | 70.726.68 | 71.606.99 | 49.774.79 | 75.172.58 | 42.117.31 | |
| RP | 87.644.26 | 81.505.36 | 89.994.00 | 81.294.96 | 86.554.09 | 91.824.34 | 94.801.01 | 85.331.63 | 47.557.02 | |

### v2

# Roundtable Policy: Confidence-Weighted-Consensus Aggregation Improves Multi-Agent-System Reasoning

@@ Abstract
Multi-agent systems have demonstrated exceptional performance in downstream tasks beyond diverse single agent baselines. A growing body of work has explored ways to improve their reasoning and collaboration, from vote, debate, to complex interaction protocols. However, it still remains opaque why specific choice would be preferred in multi-agent systems. Inspired by the decision-making mechanism of democratic committees and The Society of Mind, we introduce Roundtable Policy, an inference-time reasoning framework for multi-agent systems that performs inference through the weighted consensus of multiple LLMs. Through extensive experiments, we demonstrate its that this approach significantly enhances reasoning in complex heterogeneous scientific tasks. Roundtable Policy emphasizes structured and interpretable inference rather than opaque convergence, while requires only black-box access and uniform procedures, making it broadly applicable to diverse multi-agent systems.
@@ 1 Introduction
Scientific reasoning often requires synthesizing multiple, partially reliable perspectives. This process needs not only predicting a single correct answer, but also forming a coherent and reliable consensus across partially correct, incomplete, or even conflicting explanations. Large language models (LLMs) have shown remarkable capabilities in generating informative responses, motivating their use in scientific discovery and reasoning tasks. However, when applied to answering complex scientific problems, individual models may hallucinate facts, make brittle logical jumps, or emphasize different aspects of a problem.
These behaviors reflect not only limitations of individual models, but also a broader challenge: scientific reasoning is not fully captured by isolated, single-shot predictions, but instead requires mechanisms for aggregating multiple uncertain and heterogeneous reasoning paths. For instance, to write a scientific proposal must not only provide factually correct content, but also organize it into a structured and persuasive narrative. Addressing these requirements calls for mechanisms that can reconcile multiple reasoning paths.
@@
A natural motivation is therefore to leverage multi-agent systems, collecting their complementary strengths via diverse reasoning trajectories. Prior work has explored from majority (Wang et al., 2023), debate (Du et al., 2024; Liang et al., 2024), to other ensemble approaches (Chen et al., 2024; Wang et al., 2025). While effective in many cases, these approaches exhibit inherent limitations: majority vote, treat all models equally, overlooking that different models excel in different domains; Debate-style methods, while sometimes converging to better outputs, remain inherently opaque: it is unclear why a particular answer was selected, or hard to quantify which models were more trustworthy in the process. These observations all point to a central question: how can we effectively combine complementary strengths across models to solve complex scientific tasks?
From this perspective, multi-agent reasoning can be viewed as a problem of consensus formation across heterogeneous reasoning trajectories. We instantiate this principle with Roundtable Policy, an inference-time reasoning framework that explicitly models agent reliability with uncertainty across task-specific dimensions (Figure 2), producing auditable signals to make inference more transparent and interpretable without modifying base models. In our framework, diverse agents first propose candidate responses, which are evaluated and recorded across tasks. Then these historical evaluations are distilled into a confidence-weight table that captures both agent reliability and uncertainty. During inference, this structured memory guides aggregation, enabling the system to integrate diverse perspectives into a coherent consensus while remaining transparent about how trust is allocated across agents.
This perspective also motivates how scientific reasoning systems should be evaluated. Current evaluation has largely relied on benchmarks with isolated, singe-domain tasks. Representative examples such as GSM8K (Cobbe et al., 2021), PubMedQA (Jin et al., 2019), and AlphaGeometry (Trinh et al., 2024) have been instrumental in assessing domain-specific knowledge and pointwise accuracy. However, these benchmarks typically assume short-context, single-problem settings, and therefore fail to capture key characteristics of scientific contexts, which often require integrating knowledge across domains and maintaining logical consistency over long contexts. For example, writing a scientific proposal involves synthesizing background knowledge, justifying methodologies, and keeping long-context consistency, rather than solving an isolated arithmetic problem like 4+23*6+24-24*12 (GSM8K). This mismatch becomes sharper in multi-agent settings, where reasoning quality also depends on how complementary strengths are aggregated.
To address this gap, we design two complementary benchmarks that capture distinct but essential aspects of scientific reasoning: ScienceEval and ScienceNarrative. ScienceEval evaluates cross-domain reasoning by jointly querying agents on heterogeneous problems spanning geoscience (Deng et al., 2024), biology (Jin et al., 2019), mathematics (Cobbe et al., 2021), and multi-faceted physical science (§C). ScienceNarrative focuses on long-context reasoning consistency by requiring agents to generate structured scientific proposals under contextual templates. Across both benchmarks, Roundtable Policy consistently improves factual reliability and narrative coherence, achieving average gains of 13.01% in ScienceEval and 11.04% in ScienceNarrative over diverse baseline agents.
In summary, our contributions are threefold. (1) We formulate multi-agent reasoning as a problem of consensus formation across heterogeneous reasoning trajectories, and instantiate this perspective with Roundtable Policy, an inference-time reasoning framework. (2) We introduce two complementary benchmarks, ScienceEval and ScienceNarrative, to evaluate cross-domain and long-context consistency and accuracy, providing domain-authentic testbeds aligned with the demands of complex scientific reasoning. (3) We demonstrate the efficacy of Roundtable Policy through extensive experiments.
@@ 2 Methodology
In this section, we begin with the intuition behind the Roundtable Policy and then describe the operational mechanism of this framework.
@@ 2.1 Design Intuitions behind Roundtable Policy
In multi-agent reasoning, agents often differ in expertise and reliability across tasks. Effective collaboration therefore requires inference-time mechanisms that can account for these differences, rather than treating all agents or reasoning paths as equally trustworthy. A useful intuition comes from democratic committees. Members independently evaluate a proposal, but final decisions are rarely made by simple voting. Instead, assessments are weighted by expertise, past reliability, and confidence, leading to decisions that are both robust and interpretable. Roundtable Policy follows this intuition by viewing multi-agent reasoning as a committee-style inference process: agents propose diverse reasoning trajectories, their task-specific reliability is assessed over time, and a final consensus is formed by weighting historical reliability and confidence of each member.
2.2 Multi-Phase Setting
To realize this intuition, we structure multi-agent reasoning into multiple phases that explicitly separate proposal, evaluation, and consensus formation. Notations follow §A.
Proposition Phase.
In the proposition phase, a set of player (or proposer) agents() independently generate candidate responses to the same query. This design encourages diversity in reasoning paths and solution strategies, providing a rich pool of alternative perspectives for subsequent aggregation.
Inference Phase.
In the inference phase, an aggregator (or fusion) agent() synthesizes the candidate responses produced by the players. The aggregation is explicitly conditioned on each player’s historical performance profile, represented as a score with associated uncertainty. By incorporating this structured reliability information, the aggregator can weigh contributions beyond instantaneous agreement, enabling informed consensus formation.
Backward Phase.
In the backward phase, a committee of LLM-based grader agents() evaluates the players’ responses under task-specific or rubric-based criteria. These evaluations yield quantitative assessments of both quality and uncertainty, which are used to update the historical performance records of each player. Importantly, this phase updates only the external confidence-weight table and does not modify the parameters of any underlying language model.
2.3 Encoding Historical Performance via Confidence-Weight Table
@@
In the backward phase, each evaluation yields a trust weight with two components: (i) a quality score in , where negative values indicate incorrect or misleading responses, and (ii) a calibrated 95% confidence interval representing the uncertainty. Together, these form the equivalent of both a “grade” and a measure of “grading confidence,” ensuring that downstream decisions reflect not only evaluation outcomes but also their reliability.
@@
At round , a query is uniformly sampled from the task dataset. The -th player generate responses , which are then concatenated into . Graders then evaluate these responses, producing , which are averaged across graders to yield . After rounds, the cumulative evaluations are distilled into a confidence-weight table:
@@
The confidence-weight table (index represents agent dimension, index represents task dimension)serves as a structured memory of each player’s historical reliability across subtasks (ScienceEval) or rubric dimensions (ScienceNarrative). Unlike fine-tuning approaches, Roundtable Policy does not alter the parameters of the underlying LLMs. Instead, is iteratively updated from repeated rounds of evaluation, much like how a committee gradually learns which members are most reliable in which topics.
@@ 2.4 Inference with the Confidence-Weight Table
Once has been learned, it guides consensus aggregation at the inference phase. Given a new query , each player generates a response . The collection is passed to the fusion agent , which aggregates them using the pre-constructed :
The fusion process itself is not fine-tuned; it acts as an orchestrator that conditions on to produce a consensus answer. This mimics how a committee chair integrates both the content and credibility of reviewers before making a final decision.
@@ 3 Experiments
We design extensive experiments to evaluate the inference-time reasoning framework Roundtable Policy. Our analysis is guided by the following research questions:
RQ1(Efficacy): To what extent does Roundtable Policy enhance reasoning compared with baseline models in complex scientific environments? (see §3.2)
RQ2 (Stability & Robustness): How stable and robust is Roundtable Policy across diverse hyperparameter settings and design choices, and does it consistently provide reliable performance improvements? (see §3.3)
RQ3(Multi-Agent Spectrum): Does Roundtable Policy beat other multi-agent baselines? (see §3.4)
RQ4(Grading Rationality): To what degree do AI graders introduce bias, and can the committee setting mitigate these biases to achieve greater inter-grader consistency? (see §3.5)
3.1 Experimental Setup
Datasets & Evaluation Metrics. To simulate realistic scientific reasoning environments, we consider two complementary evaluation regimes that reflect distinct but equally important aspects of scientific work: (i)ScienceEval: cross-domain, heterogeneous problem solving (Figure 4) (ii)ScienceNarrative: coherent, narrative-driven scientific writing (Figure 3). These settings are together designed to capture the dual demands of scientific intelligence—answering pointwise questions and constructing structured, logically consistent arguments. Following is details for both settings:
| Setting | Model | Multi-Faceted Physical Reasoning | Other Domains | |||||||
@@ 3.1 Experimental Setup
| ScienceEval |
GPT-4o
85.894.27 | 78.265.37 | 87.714.11 | 74.595.49 | 85.144.40 | 85.584.60 | 87.812.13 | 83.552.20 | 51.536.64 |
Grok-2
83.924.07 | 76.244.96 | 85.613.92 | 76.184.99 | 82.024.10 | 82.634.39 | 88.841.70 | 77.392.28 | 51.316.25 | |
Claude-3.5H
82.074.65 | 71.206.04 | 84.724.37 | 71.495.83 | 77.745.03 | 77.955.22 | 81.192.58 | 73.682.58 | 48.526.75 | |
GPT-4.5p
72.866.77 | 71.466.96 | 76.676.12 | 72.786.36 | 77.185.71 | 77.686.17 | 82.502.68 | 82.592.08 | 50.636.89 | |
DeepSeek-chat
77.635.35 | 72.026.14 | 81.534.97 | 72.965.79 | 79.644.87 | 79.815.23 | 72.233.16 | 78.052.41 | 47.836.57 | |
Claude-3.7S
77.433.46 | 70.514.05 | 79.243.13 | 72.393.70 | 77.413.49 | 79.493.12 | 62.342.93 | 72.561.97 | 47.765.36 | |
Llama-3.3
72.145.96 | 63.866.95 | 75.335.55 | 67.136.30 | 71.235.71 | 73.875.72 | 87.431.93 | 72.772.49 | 48.646.54 | |
Qwen-Plus
80.325.01 | 72.366.03 | 82.974.71 | 73.425.66 | 81.494.70 | 81.184.93 | 86.992.08 | 76.272.48 | 51.646.57 | |
Gemini-2.0
67.267.61 | 61.618.09 | 74.966.58 | 65.737.31 | 70.726.68 | 71.606.99 | 49.774.79 | 75.172.58 | 42.117.31 | |
| Ours(RP) | 87.644.26 | 81.505.36 | 89.994.00 | 81.294.96 | 86.554.09 | 91.824.34 | 94.801.01 | 85.331.63 | 47.557.02 | |